40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jul 30 11:05:44 2018, gautam, Update, SUS, IFO SUS wonky IMG_7102.JPG
    Reply  Mon Jul 30 16:11:54 2018, gautam, Update, SUS, Trillium interface box is broken 
       Reply  Mon Jul 30 18:19:03 2018, Koji, Update, SUS, Trillium interface box was fixed and reinstalled P_20180730_173134.jpgP_20180730_180151.jpg
          Reply  Tue Jul 31 08:17:55 2018, Steve, Update, SUS, Trillium interface box was fixed,reinstalled & working all_OK.png
Message ID: 14115     Entry time: Mon Jul 30 11:05:44 2018     Reply to this: 14117
Author: gautam 
Type: Update 
Category: SUS 
Subject: IFO SUS wonky 

When I came in this morning:

  • PMC was unlocked.
  • Seis BLRMS were off scale.
  • ITMX OSEM LEDs were dark on the CRT monitor even though Sat Box was plugged in.

Checking status of slow machines, it looked like c1sus, c1aux, and c1iscaux needed reboots, which I did. Still PMC would not lock. So I did a burtrestore, and then PMC was locked. But there seemed to be waaaaay to much motion of MCREFL, so I checked the suspension. The shadow sensor EPICS channels are reporting ~10,000 cts, while they used to be ~1000cts. No unusual red flags on the CDS side. Everything looked nominal when I briefly came in at 6:30pm PT yesterday, not sure if anything was done with the IFO last night.

Pending further investigation, I'm leaving all watchdogs shutdown and the PSL shutter closed.

A quick look at the Sorensens in 1X6 revealed that the +/- 20V DC power supplies were current overloaded (see Attachment #1). So I set those two units to zero until we figure out what's going on. Possibly something is shorted inside the ITMX satellite box and a fuse is blown somewhere. I'll look into it more once Steve is back.

Attachment 1: IMG_7102.JPG  1.349 MB  Uploaded Mon Jul 30 12:31:35 2018  | Hide | Hide all
IMG_7102.JPG
ELOG V3.1.3-